If the temperature of the sun is estimated to be 5700k, calculate the total power it radiates.Assume the sun is a spherical blackbody of radius 7.1*10^5km
1
Expert's answer
2020-08-10T19:54:54-0400

According to Stefan-Boltzmann law, the luminosity (the power of radiation) of the object can be calculated as

"L=4\pi R^2\sigma T^4 = 4\pi \cdot (7.1\cdot10^8)^2\cdot 5.67\cdot10^{-8}\cdot (5700)^4 = 3.8\cdot10^{26}\,\mathrm{W}."